# Optimal parallelisation strategies for flat histogram Monte Carlo sampling

### v2

@@ III.5 Results summary
It can be seen that all of the proposed methods achieve an appreciable speedup, with even the worst-performing methods (5 and 6) achieving near linear speedup out to 16 WL instances, i.e. energy sub-domains. However, the methods which stand out are 1, 2, 3, and 4, i.e. those using non-uniform energy domain decomposition. This result reinforces the idea that the greatest gain in performance is obtained via non-uniform domain decomposition. Methods 1 and 2, which make use of the proposed dynamic energy sub-domain sizing scheme, offer modest improvements over the case of statically-sized non-uniform energy sub-domains. The increase in performance between the load-balanced (Methods 1, 2) and not load balanced (Methods 3, 4) allows for the simulation to retain its sub-domain efficiency for a greater number of sub-domains giving access to greater speedups. Of note, replica exchange does not appear to significantly impact performance either positively or negatively.
@@ IV Summary, Conclusions, and Recommendations
The most significant factor in determining parallel performance is the choice of energy domain decomposition scheme, with use of non-uniform energy sub-domains consistently providing the largest gains in both speedup and overall simulation efficiency. Dynamic load balancing offers a modest further enhancement in performance by allowing the energy domain decomposition scheme to adjust itself ‘on the fly’ and adapt to the sampling demands of complex energy landscapes, with energy sub-domain boundaries shifting between each WL iteration. Careful attention to the number of walkers allocated to each energy sub-domain also proves critical. While employing multiple walkers may seem like a natural choice when seeking to parallelise the WL sampling algorithm, our results indicate that this produces rapidly diminishing returns as a function of number of walkers, and potentially a reduction in efficiency. In this study, we found that the use of one or two walkers per sub-domain is generally sufficient to realise any performance improvements available through the use of multiple walkers per energy sub-domain. Also of note is that replica exchange does not noticeably influence efficiency in this test case, but nor does it compromise performance, which suggests that it may be included where needed to improve sampling without penalty. This is likely due to the non-physical MC moves (long-range atomic swaps) in our test system which result in the ergodicity not being limited by the need to cross energy barriers. Finally, the construction of overlap regions between sub-domains is shown to have a negligible effect on performance unless it is in excess of a 75% overlap, underscoring that large overlap is necessary only for replica exchange and not for the reliable combination of local DOS histograms to form the global DOS.
@@ IV.1 Recommendations
On the basis of these findings, we recommend that parallel implementations of WL sampling prioritise using non-uniform energy domain decomposition as a primary route to improved simulation efficiency. Subsequently, one should implement multiple walkers per sub-domain for further efficiency and performance. Once those two features have been implemented, dynamic load balancing should subsequently be incorporated wherever feasible for modest improvement of per-instance efficiency. If reduced runtime is more important than efficiency, and if adding further sub-domains hurts performance, further increasing the number of walkers can be employed as a complementary tool, though they will decrease computational efficiency. Finally, overlap regions of a modest size should be included primarily to facilitate replica exchange, as they have little to no impact on the accuracy of the reconstructed total DOS, though excessively-sized overlap regions should be avoided due to the associated drop in parallel performance.
@@
Hubert J. Naguszewski Conceptualization (equal), Data curation, Investigation, Methodology (lead), Software (lead), Validation, Visualization, Writing - original draft (lead), Writing - review and editing (equal). Christopher D. Woodgate Methodology (supporting), Software (supporting), Supervision (supporting), Writing – original draft (supporting), Writing – review and editing (equal). David Quigley Conceptualization (equal), Funding acquisition, Project administration, Resources, Supervision (lead), Writing – review and editing (equal).
@@
The data supporting the findings of this study can be made available upon request. The BraWl package, which was used to obtain all results presented in this work is fully open-source and available at https://github.com/ChrisWoodgate/BraWl.

### v3

@@ III.1.5 AlTiCrMo results summary
Focusing on plot (a) of Figure 10, it can be seen that all of the proposed methods achieve an appreciable speedup, with even the worst-performing methods (5 and 6) achieving near linear speedup out to 16 WL instances, i.e. energy sub-domains. However, the methods which stand out are 1, 2, 3, and 4, i.e. those using non-uniform energy domain decomposition. This result reinforces the idea that the greatest gain in performance is obtained via non-uniform domain decomposition. Methods 1 and 2, which make use of the proposed dynamic energy sub-domain sizing scheme, offer modest improvements over the case of statically-sized non-uniform energy sub-domains. The increase in performance between the load-balanced (Methods 1, 2) and not load balanced (Methods 3, 4) allows for the simulation to retain its sub-domain efficiency for a greater number of sub-domains giving access to greater speedups. Of note, replica exchange does not appear to significantly impact performance either positively or negatively.
III.2 CuZn
Now that the effects of the parallelisation schemes have been outlined for AlTiCrMo, we will contrast the results to the more simple CuZn alloy. For this benchmark, we use a simulation supercell made up of $8\times 8\times 8$ bcc cubic unit cells for a total of 1024 atoms. We find that 1024 atoms is necessary for an appreciable number of energy bins for the purposes of load balancing due to the degenerate energy landscape found within this binary alloy. When running the same number of atoms as with the AlTiCrMo case (128 atoms) the simulation only requires 16 bins and takes $\sim$2.5 seconds to complete 12 WL iterations (one WL instance across the entire energy domain), further demonstrating the challenging nature of the AlTiCrMo system. Parallelisation is, however, still useful in the simpler CuZn case when simulating a larger system. Our WL sampling simulations used $N=256$ energy bins over an energy domain of E=76E=-76 meV/atom to E=5E=-5 meV/atom, chosen via the method outlined in Sec. II.1 with $T_{\text{min}}=50K$ and $T_{\text{max}}=3000K$.
For CuZn, we select the two most informative metrics for comparison: the relative MC steps and total speedup. These results can be seen in plots (b) of Figures 9 and 10 respectively. Starting with plot (b) in Figure 9, it can be seen that the total number of MC steps required to converge the simulation drops once non-uniform energy sub-domains are introduced. This effect is less pronounced than the AlTiCrMo case, however panels (b.i) and (b.ii) display a similar trend of an initial reduction in MC steps required before subsequently increasing. Panel (b.iii) is of particular interest, as it exhibits an evidently different behaviour, with increasing the number of uniform domains directly increasing the total number of MC steps required to converge the simulation. This increase in the required work, which would also apply to the methods in panels (b.i) and (b.ii), is caused by the difficulty of converging low energy bins. Using the case in panel (b.iii), as more sub-domains are introduced their size decreases, however, the difficult to converge sub-domain never becomes small enough to effectively divide the load. This issue appears in cases (b.i) and (b.ii) due to the limit imposed on the minimum size of a sub-domain described in Section II.2.4. In both cases this results in an overall increase in the number of MC steps, as each iteration requires the low energy domain to converge before completing.
Proceeding, we now look to Figure 10 and begin by addressing the difference in magnitude of speedup between AlTiCrMo and CuZn. CuZn exhibits a decreased overall benefit from parallelisation, which is most evident by the speedup not reaching the theoretical $h^{2}$ threshold. This can be attributed to the discretised DOS and degenerate energy states (as compared to AlTiCrMo) which result in particularly non-uniform convergence times across the energy domain. Both plots also exhibit a similar trend between methods that have and do not have non-uniform domain decomposition, which shows that the key to unlocking higher parallelisation is the introduction of non-uniform domains (particularly evident in the CuZn case).
@@ IV Summary, Conclusions, and Recommendations
The most significant factor in determining parallel performance is the choice of energy domain decomposition scheme, with use of non-uniform energy sub-domains consistently providing the largest gains in both speedup and overall simulation efficiency. Dynamic load balancing offers a modest further enhancement in performance by allowing the energy domain decomposition scheme to adjust itself ‘on the fly’ and adapt to the sampling demands of complex energy landscapes, with energy sub-domain boundaries shifting between each WL iteration. Careful attention to the number of walkers allocated to each energy sub-domain also proves critical. While employing multiple walkers may seem like a natural choice when seeking to parallelise the WL sampling algorithm, our results indicate that this produces rapidly diminishing returns as a function of number of walkers, and potentially a reduction in efficiency. In this study, we found that the use of one or two walkers per sub-domain is generally sufficient to realise any performance improvements available through the use of multiple walkers per energy sub-domain. Also of note is that replica exchange does not noticeably influence efficiency in this test case, but nor does it compromise performance, which suggests that it may be included where needed to improve sampling without penalty. This is likely due to the non-physical MC moves (long-range atomic swaps) in our test system which result in the ergodicity not being limited by the need to cross energy barriers. Finally, the construction of overlap regions between sub-domains is shown to have a negligible effect on performance unless it is in excess of a 75% overlap, underscoring that large overlap is necessary only for replica exchange and not for the reliable combination of local DOS histograms to form the global DOS. We would also like to highlight that the biggest overall reductions in the total MC steps needed to converge the WL sampling simulations are more prominent for the more challenging of the two benchmark systems (AlTiCrMo) and we expect the largest benefit from the parallelisation schemes to be found in such scenarios, i.e. complex systems with quasi-continuous energy landscapes.
@@ IV.1 Recommendations
On the basis of these findings, we recommend that parallel implementations of WL sampling prioritise using non-uniform energy domain decomposition as a primary route to improved simulation efficiency. Subsequently, one should implement multiple walkers per sub-domain for further efficiency and performance. Once those two features have been implemented, dynamic load balancing should subsequently be incorporated wherever feasible for modest improvement of per-instance efficiency. If reduced runtime is more important than efficiency, and if adding further sub-domains hurts performance, further increasing the number of walkers can be employed as a complementary tool, though they will likely decrease overall computational efficiency. Finally, overlap regions of a modest size should be included primarily to facilitate replica exchange, as they have little to no impact on the accuracy of the reconstructed total DOS, though excessively-sized overlap regions should be avoided due to the associated drop in parallel performance.
@@
Hubert J. Naguszewski Conceptualization (equal), Data curation, Investigation (lead), Methodology (lead), Software (lead), Validation, Visualization, Writing - original draft (lead), Writing - review and editing (equal). Christopher D. Woodgate Investigation (supporting), Methodology (supporting), Software (supporting), Supervision (supporting), Writing – original draft (supporting), Writing – review and editing (equal). David Quigley Conceptualization (equal), Funding acquisition, Project administration, Resources, Supervision (lead), Writing – review and editing (equal).
@@
The data supporting the findings of this study are freely available at the following DOI/URL: https://doi.org/10.5281/zenodo.18921874. The Brawl package, which was used to obtain all results presented in this work is fully open-source and available at https://github.com/ChrisWoodgate/BraWl.